Restart Night #3 UT 2018-09-04

OSA: S. Allanson, J. Williams

Observers/ISA: B. Rothberg, J. Hill (Tucson)

ISp: E. Solheid

Mtn Manager: P. Hartley


UT 00:29 - precipation at summit

UT 01:39 - Handing off to Dave T for LUCI2 wavelength calibrations

UT 03:09 - Powered up LBCs (had to restart once) - taking test Biases

UT 03:15 - LBC-Red filter error, biases stopped at 7/10. Trying again

Open

UT 04:00 Steve and Josh have completed eyeball calibration of SX, and are now adjusting AGW4 guiding hotspot.

UT 04:12 Hotspot update seems to block WFS loop, but not even a pause indication. Is this seeing + hotspot change > 0.5 sec? Why doesn't it recover?

2018-09-04T04:11:20.807869+00:00 tcs1 LBT_GCSL: GT: WFS exposure (9.0s) still blocked.  PR2HS (0.991) > maxHotspotFromProbe , -0.991 < 0.500 NS Rates: -0.071 -0.023 [mm/hr], Tot: 0.124[as/hr], PR2HS_Sign: -1, KFP2ProbeTime: 5.408 [s], KFP2ProbeTimeAvg: 5.653 [s]
2018-09-04T04:11:29.522583+00:00 tcs1 LBT_GCSL: GT: WFS exposure (9.0s) still blocked.  PR2HS (0.991) > maxHotspotFromProbe , -0.990 < 0.500 NS Rates: -0.079 -0.009 [mm/hr], Tot: 0.133[as/hr], PR2HS_Sign: -1, KFP2ProbeTime: 5.513 [s], KFP2ProbeTimeAvg: 5.611 [s]

UT 04:13 New preset

UT 04:26 stuck paused at 250,250. edit LUCI_L.cfg and send new preset, but need to restart GCS to reread/capture the saved hotspot.

2018-09-04T04:25:42.628017+00:00 tcs1 LBT_GCSL: GT: warning WFS exposure in progress, but moving the probe, pause WFS
2018-09-04T04:25:42.642804+00:00 tcs1 LBT_GCSL: WFS: pauseLoop: pausing WFS loop.
2018-09-04T04:25:42.671825+00:00 tcs1 LBT_GCSL: WFS: triggerWFS: abort waiting for image as a pause request was received..
2018-09-04T04:25:42.701002+00:00 tcs1 LBT_GCSL: WFS: committing WFS telemetry sample
2018-09-04T04:25:50.607669+00:00 tcs1 LBT_GCSL: GT: block next WFS exposure (4.0s), 0.948 >= 0.500 NS Rates: -0.075 -0.004 [mm/hr], Tot: 0.125[as/hr], PR2HS_Sign: 1, PR2HS: 0.947, KFP2ProbeTime: 5.341 [s], KFP2ProbeTimeAvg: 5.393 [s]
2018-09-04T04:25:59.085770+00:00 tcs1 LBT_GCSL: GT: WFS exposure (4.0s) still blocked.  PR2HS (0.947) > maxHotspotFromProbe , -0.947 < 0.500 NS Rates: -0.066 -0.014 [mm/hr], Tot: 0.112[as/hr], PR2HS_Sign: -1, KFP2ProbeTime: 5.267 [s], KFP2ProbeTimeAvg: 5.355 [s]
2018-09-04T04:26:08.338809+00:00 tcs1 LBT_GCSL: GT: WFS exposure (4.0s) still blocked.  PR2HS (0.947) > maxHotspotFromProbe , -0.947 < 0.500 NS Rates: -0.065 -0.008 [mm/hr], Tot: 0.108[as/hr], PR2HS_Sign: -1, KFP2ProbeTime: 5.803 [s], KFP2ProbeTimeAvg: 5.489 [s]
2018-09-04T04:26:14.729834+00:00 tcs1 LBT_GCSL: WFSCam: temperatures ccd: -999.9 C, dewar: -999.9 C
2018-09-04T04:26:14.729848+00:00 tcs1 LBT_GCSL: [D] WFS: readCamTemp OK.
2018-09-04T04:26:17.421808+00:00 tcs1 LBT_GCSL: GT: WFS exposure (4.0s) still blocked.  PR2HS (0.947) > maxHotspotFromProbe , -0.946 < 0.500 NS Rates: -0.027 -0.132 [mm/hr], Tot: 0.225[as/hr], PR2HS_Sign: -1, KFP2ProbeTime: 5.870 [s], KFP2ProbeTimeAvg: 5.604 [s]

04:32 Changed max hotspot distance from 0.54 arcsec to 2.54 arcsec in LUCI_L.cfg (TEMPORARY), but doesn't take effect until new preset. I changed maxOffsetToHotspot, but perhaps maxHotspotFromProbe is also relevant.

-Y moves the spots up.

04:38 246,246

04:44 250,246

04:47 256,246 release Zernikes.

04:54 restart GCSL with hotspot 264,246, resend preset.

05:01 266,246

Pupil center measured by IDL 159.4, 218.0 this made both GCSL and IDL agree, when they had opposite signs a few minutes ago. WHY???

Releasing Zernikes has Z5 running away. Called Doug to remember how we changed this. Petr changed this in GCS to flip the raw Zernikes and Doug has an if statement in IDL.

05:32 enabled flipX in GCS LUCI_L.cfg, and Doug is adding AGW4 to his IDL if statement

changed sext_x to match hotspot value (just to be sure).

05:40 changed lenslet edge from 174.7 to 172.7 and blew up GCS collimation. Steve and Josh recovered very nice collimation by eye.

Doug says I should be able to change lenslet_edge in LUCI_L.cfg to make "Pupil Center" match the IDL value.

gcs_wfsc_display, /show_pupil

Then use the "Outer Pupil" position that IDL prints.

06:04 Set pupil position to 170.7,218.6 in both IDL and GCS. (didn't take effect until 06:21 UT when we sent the new preset)

06:10 redoing pointing to get back to the correct guide star. SX IE=-63 CA=+13 DX IE=-135 CA=+20

06:20 ACTIVE preset to BS9108 ?

06:28 restarting GCSL still doesn't change the GCS value of Pupil Center. We are trying to understand why. GCS doesn't print the "lenslet edge" value.

06:43 Outer Pupil Fit is 166.17,218.58 slightly different than the 170.7,218.6 that Doug measured with ds9 earlier.

06:47 lenslet edge 180.7,226.7 (from 174.7,226.7)

06:49 send new preset to GCSL

06:52 stop/start GCSL and send new preset.

06:59 set lenslet edge to -1, and send new preset

07:02 edit in new lenslet_edge lines from LUCI_R (to avoid non-printing characters)

07:04 stopt/start GCSL and send new preset.

07:07 stopWFSing, and collimate_once with IDL = collimate_once, /send_zern IDL converged OK, although it took a few iterations to start correcting Z22. Image is a bit triangular in 1 arcsec seeing.

07:25 change guiding hotspot 267.0.246.0 (from 270,246) (do this with GCSGUI Window while collimating with IDL) That seems to be the wrong way, so change to 273,246.

See Issue 7388 about this lenslet_edge problem.

X,Y average removed was +0.3 +0.1 in units of fractional lenslets. 273,247 is the new hotspot. It's way easier to do this while collimating with IDL.

SX seems to have funny image jumps - like briefly out of focus every once in a while.

UT 07:40 Switch over to LBC.

UT 08:14 - sending co-pointing script for Little Dumbbell - co-pointing limit reached - error

UT 08:21 - used old CP script for IC 342 (low elevation) that worked, the CP script generated by the OT violated the co-pointing limit by immediately sending an asynch error

-hypothesis is that the units of proper motion from the OT script (generated on the backend) is radians/century not "/year

UT 08:32 - running co-pointing (airmass 1.46)

UT 08:40 - co-pointing done, co-pointing was ok, pointing was bad

UT 08:41 - running PA=0 superfoc sequence (airmass 1.44)

UT 08:44 - preset cancelled just before running superfoc -so ignore data

UT 08:45 - resent CP preset IC 342 - superfoc PA=0 84644 (blue) 84636 (red)

UT 08:49 - DIMM 0.75"

UT 08:52 - running PA=180 superfoc 85355 (blue) 85353 (red)

UT 09:00 - running DOFPIA DIMM 0.64"

UT 09:02 - Running PA=0 rotator center script

UT 09:05 - Running PA=180 rotator center script

UT 09:07 - Running PA=90 rotator center script

UT 09:10 - Running PA=270 rotator center script last files 91202 (blue) 91157 (red)

UT 09:14 - Testing hypothesis - re-running IC342 CP with Proper Motions = 0

UT 09:15 - Running DOFPIA DIMM 0.73"

UT 09:20 - Sending IC 342 science OB - testing dither and guiding (airmass 1.36)

UT 09:31 ~0.75" LBC-Blue, ~0.6" LBC-Red

UT 09:53 - stopping IC342 after 3 dithers (plus 2 images in blue, red at 4th dither position)

UT 09:54 - CP script for Little Dumbbell Nebula (collimate here)

UT 09:57 - selected the wrong script - RA=19hr - RIP'd the shells, OSA stopped at elevation 18 - recovering

UT 10:00 - resending Little Dumbbell Nebula CP script

UT 10:02 - running DOFPIA

UT 10:08 - running CP script for Little Dumbbell

UT 10:12 - Running superfocus (airmass 1.05) PA=0 101239 (blue) 101237 (red)

UT 10:18 - Running superfocus (airmass 1.05) PA=180 101938 (blue) 101932 (red)

UT 10:25 - running DOFPIA - DIMM 0.84"

UT 10:32 - running M76 BI script

UT 10:47 - running DOFPIA -DIMM 0.77"

UT 10:51 - running M76 UR script

UT 11:07 - moving to NS test of 2017WV13 CP field

UT 11:10 -Running DOFPIA

UT 11:14 - Co-pointing for 2017WV13

UT 11:19 - NSIGUI override - sending script for 2017WV13 5pt dither (used the UT 11 coords)

UT 11:34 - Running DOFPIA - but need to kill NS preset first -oops

UT 11:35 -re-Running DOFPIA (with CP script)

UT 11:42 - NSIGUI override - sending script for 2017WV13 5pt dither (used the UT 12 coords)

UT 11:58 - finished scripts

UT 12:00 - powering off LBCs

UT 12:01 - running allseeing


LBCR rotator center:
lbcr.20180904.090355.fits any 0.16874
lbcr.20180904.090619.fits any -179.72181
lbcr.20180904.090905.fits any 90.38879
lbcr.20180904.091157.fits any -89.71879

090355 & 090619 -> (1080.46,2913.19)
090905 & 091157 -> (1078.44, 2911.99)
090355 & 090905 -> (1081.72, 2917.22)
090619 & 091157 -> (1076.00,2909.97)

Adopt average of results from diametrically opposed pairs: ~(1079.5,2912.5)

LBCB rotator center:
lbcb.20180904.090358.fits any 0.16666
lbcb.20180904.090626.fits any -179.72576
lbcb.20180904.090911.fits any 90.39081
lbcb.20180904.091202.fits any -89.72048

090358 & 090626 -> (1033.27,2917.94)
090911 & 091202 -> (1032.5, 2919.5)
090358 & 090911 -> (1036.1,2922.73)
090626 & 091202 -> (1036.98,2922.77) (few stars found)

Adopt average of first two: ~(1033,2919)

In headers as CRPIX1, CRPIX2 for chip 2 are:
Red: 1078, 2913
Blue: 1035, 2924

Superfoc sequences:

Best-focus at center of chip2 - values appear to be reasonable but will be listed here when ready.

Chip flatness: TBD. gnuplot 3D script not rotating on mac.
Topic revision: r11 - 07 Sep 2018, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback